# LaFiCMIL: Rethinking Large File Classification from the Perspective of Correlated Multiple Instance Learning

### v3

@@ Abstract
Transfomer-based models have significantly advanced natural language processing, in particular the performance in text classification tasks. Nevertheless, these models face challenges in processing large files, primarily due to their input constraints, which are generally restricted to hundreds or thousands of tokens. Attempts to address this issue in existing models usually consist in extracting only a fraction of the essential information from lengthy inputs, while often incurring high computational costs due to their complex architectures. In this work, we address the challenge of classifying large files from the perspective of correlated multiple instance learning. We introduce LaFiCMIL, a method specifically designed for large file classification. LaFiCMIL is optimized for efficient operation on a single GPU, making it a versatile solution for binary, multi-class, and multi-label classification tasks. We conducted extensive experiments using seven diverse and comprehensive benchmark datasets to assess LaFiCMIL’s effectiveness. By integrating BERT for feature extraction, LaFiCMIL demonstrates exceptional performance, setting new benchmarks across all datasets. A notable achievement of our approach is its ability to scale BERT to handle nearly 20 0002000020\,00020 000 tokens while operating on a single GPU with 32GB of memory. This efficiency, coupled with its state-of-the-art performance, highlights LaFiCMIL’s potential as a groundbreaking approach in the field of large file classification.
LaFiCMIL: Rethinking Large File Classification from the Perspective of Correlated Multiple Instance Learning
@@ 1 Introduction
Text classification is a fundamental task in Natural Language Processing (NLP), entailing the assignment of suitable label(s) to specific input texts Kowsari et al. (2019); Premasiri et al. (2023). This process is crucial across various domains, including sentiment analysis Dang et al. (2020), fake news detection Kumar et al. (2020), and offensive language identification Ranasinghe and Zampieri (2020), among others. Recent years have seen the emergence of self-attention-based models like Transformer Vaswani et al. (2017), GPT Radford et al. (2018, 2019), and the BERT family Devlin et al. (2018); Feng et al. (2020); Sun et al. (2023), which have established state-of-the-art benchmarks in text classification tasks. However, the challenge of processing very long documents remains a significant obstacle, largely due to the high computational requirements of these models when facing extremely large number of tokens.
There are mainly two types of solutions in the literature to address long token sequences: ① extending the input length limit using a sparse attention mechanism, such as Longformer Beltagy et al. (2020), and ② dividing long documents into segments and recurrently processing the Transformer-based segment representations, such as RMT Bulatov et al. (2022, 2023). Nevertheless, they either struggle with handling extremely long sequences, as with Longformer, or suffer from information loss across the recurrent processing of segments, as with RMT.
In this work, we leverage Multiple Instance Learning (MIL) to tackle the problem of large file classification. MIL deals with a bag of instances for which only a single bag-level label is assigned, while instance-level labels remain unknown. Furthermore, the number of instances in each bag is undetermined, necessitating a flexible MIL model capable of accommodating input bags with varying instance counts. Recently, MIL has been successfully applied to computer vision problems, particularly whole slide image classification Shao et al. (2021); Zhang et al. (2022). This inspired us to leverage MIL to address the large file classification problem.
@@
We introduce LaFiCMIL, a simple yet effective Large File Classification approach based on correlated Multiple Instance Learning. On the one hand, as proven in Theorem 1 (cf., Section 3), a MIL score function for a bag classification task can be approximated by a series of sub-functions of the instances. This inspires us to split a large file into smaller chunks and extract their features separately using BERT. On the other hand, we aim to guide the model to learn high-level overall features from all instances, rather than deriving the final bag prediction from instance predictions based on a simplistic learned projection matrix. In addition, in contrast to the basic version of MIL Ilse et al. (2018), where instances within the same bag exhibit neither dependency nor ordering among one another, we claim that the small chunks from the same large file are correlated in some way (e.g., semantic dependencies in paragraphs). This implies that the presence or absence of a positive instance in a bag can be influenced by the other instances contained within the same bag. As a result, relying on our computationally efficient LaFiAttention layer, our approach is capable of efficiently extracting correlations among all chunks as additional information to boost classification performance.
@@
  • The training of LaFiCMIL is super efficient, which requires only 1.86$\times$ training time than the original BERT, but is able to handle 39$\times$ longer sequence on a single GPU.
@@ 2.1 Large File Classification
In recent years, significant efforts have been made to alleviate the input limit of Transformer-based models to handle different types of large files. One notable example is Longformer Beltagy et al. (2020), which extends the limit to $4096$ tokens using a sparse attention mechanism Zaheer et al. (2020). CogLTX Ding et al. (2020) chooses to identify key sentences through a trained judge model. Alternatively, ToBERT Pappagari et al. (2019) and RMT Bulatov et al. (2022, 2023) segment long documents into fragments and then aggregate or recurrently process their BERT-based representations. Recently, two simple BERT-based methods proposed in Park et al. (2022) achieved state-of-the-art performance on several datasets for long document classification. Specifically, BERT+Random selects random sentences up to 512 tokens to augment the first 512 tokens. BERT+TextRank augments the first 512 tokens with a second set of 512 tokens obtained via TextRank Mihalcea and Tarau (2004). They also provide a comprehensive evaluation to compare the relative efficacy of various baselines on diverse datasets, which revealed that no single approach consistently outperforms others across all six benchmark datasets, encompassing different classification tasks such as binary Kiesel et al. (2019), multi-class Lang (1995), and multi-label classification Bamman and Smith (2013); Chalkidis et al. (2019).
@@ 2.2 Multiple Instance Learning
Multiple Instance Learning (MIL) has attracted increasing research interests and applications in recent years. The application scenarios of MIL span across various domains Ji et al. (2020); Song et al. (2019); Hebbar et al. (2021), but the most prominent one is Medical Imaging and Diagnosis. Particularly, there has been a growing trend towards developing MIL algorithms for medical whole slide image analysis Kanavati et al. (2020); Xu et al. (2019).
These MIL models can generally be categorized into two groups based on whether the final bag predictions are derived directly from instance predictions Feng and Zhou (2017); Lerousseau et al. (2020); Lu et al. (2021a); Sharma et al. (2021) or from aggregated instance features Li et al. (2021); Shao et al. (2021); Zhang et al. (2022). For the first group, bag predictions are typically achieved through either average pooling or maximum pooling. In contrast, the second group learns a high-level representation of a bag and constructs a classifier on top of this bag representation for bag-level predictions. Although instance-level probability pooling is simple and straightforward, empirical evidence has demonstrated that it is less effective than its bag embedding counterpart Wang et al. (2018); Shao et al. (2021).
Furthermore, the fundamental assumption of Multiple Instance Learning (MIL) postulates that instances within a bag are independent of each other, a supposition that may not hold true in practical applications. Consequently, some researchers have endeavored to explore scenarios wherein instances within a bag exhibit correlations or dependencies, a concept referred to as Correlated Multiple Instance Learning (c-MIL) Zhou et al. (2009); Zhang (2021); Shao et al. (2021). This suggests that the presence or absence of a positive instance in a bag could be affected by other instances in the same bag. Nevertheless, applying c-MIL to solve large file classification problems beyond whole slide image classification remains under-explored.
@@ 3 Technical Preliminaries
In this section, we describe several essential technical preliminaries, which underpin and inform the development of LaFiCMIL. We first present a pair of theorems that substantiate the foundation of our approach, the fundamental principles of Correlated Multiple Instance Learning (c-MIL).
Theorem 1.
Suppose $S:\chi\rightarrow\mathbb{R}$ is a continuous set function w.r.t Hausdorff distance [^1] https://en.wikipedia.org/wiki/Hausdorff˙distance $d_{H}(.,.)$. $\forall\varepsilon>0$, for any invertible map $P:\chi\rightarrow\mathbb{R}^{n}$, $\exists$ function $\sigma$ and $g$, such that for any set $X\in\chi$:
$|S(X)-g(P_{X\in\chi}\{\sigma(x):x\in X\})|<\varepsilon$ (1)
The proof of Theorem 1 can be found in Shao et al. (2021). From this theorem, we can conclude that a Hausdorff continuous set function $S(X)$ can be arbitrarily approximated by a function in the form $g(P_{X\in\chi}\{\sigma(x):x\in X\})$. This insight can be applied to MIL, as the mathematical definition of sets in the theorem is equivalent to that of bags in MIL framework. Consequently, the theorem provides a foundation for approximating bag-level predictions in MIL using instance-level features.
Theorem 2.
@@
$\begin{split}H(\theta_{1},\theta_{2},...,\theta_{n})&=\sum_{t=2}^{n}H(\theta_{t}|\theta_{1},\theta_{2},...,\theta_{t-1})+H(\theta_{1})\\ &\leq\sum_{t=1}^{n}H(\theta_{t})\end{split}$ (2)
The proof of Theorem 2 can be found in Shao et al. (2021). This theorem demonstrates that the information entropy of a bag under the correlation assumption is smaller than the information entropy of a bag under the i.i.d. assumption. The lower information entropy in Correlated Multiple Instance Learning (c-MIL) suggests reduced uncertainty and the potential to provide more valuable information for bag classification tasks. In Section 4.1, we introduce c-MIL, and in Section 4.2, we derive the efficient LaFiCMIL, which learns and exploits correlations among instances to address large file classification problem.
In the remainder of this section, we present the necessary preliminaries for our efficient attention layer inspired by the Nyströmformer Xiong et al. (2021), referred to as LaFiAttention in the following discussion, which performs as a sub-function within our proposed LaFiCMIL.
In the original Transformer Vaswani et al. (2017), an input sequence of $n$ tokens of dimensions $d$, $X\in{\bf R}^{n\times d}$, is projected using three matrices $W_{Q}\in{\bf R}^{n\times d_{q}}$, $W_{K}\in{\bf R}^{n\times d_{k}}$, and $W_{V}\in{\bf R}^{n\times d_{v}}$, referred as query, key, and value respectively with $d_{k}=d_{q}$. The outputs $Q,K,V$ are calculated as
$Q=XW_{Q},\leavevmode\nobreak\ K=XW_{K},\leavevmode\nobreak\ V=XW_{V}$ (3)
@@
$D(Q,K,V)=SV=softmax(\frac{QK^{T}}{\sqrt{d_{q}}})V$ (4)
@@
$S=softmax(\frac{QK^{T}}{\sqrt{d_{q}}})=\begin{bmatrix}A_{S}&B_{S}\\ F_{S}&C_{S}\end{bmatrix}$ (5)
@@
$\hat{S}=softmax(\frac{Q\widetilde{K}^{T}}{\sqrt{d_{q}}})A^{+}_{S}softmax(\frac{\widetilde{Q}K^{T}}{\sqrt{d_{q}}}),$ (6)
where $\widetilde{Q}=[\widetilde{q_{1}};...;\widetilde{q_{m}}]\in{\bf R}^{m\times d_{q}}$ and $\widetilde{K}=[\widetilde{k_{1}};...;\widetilde{k_{m}}]\in{\bf R}^{m\times d_{q}}$ are the selected landmarks for inputs $Q=[q_{1};...;q_{n}]$ and $K=[k_{1};...;k_{n}]$, $A^{+}_{S}$ is the Moore-Penrose inverse [^2] https://en.wikipedia.org/wiki/Moore-Penrose˙inverse of $A_{S}$.
Lemma 1.
For $A_{S}\in{\bf R}^{m\times m}$, the sequence $\{Z_{j}\}^{j=\infty}_{j=0}$ generated by Razavi et al. (2014),
$Z_{j+1}=\frac{1}{4}Z_{j}(13I-A_{S}Z_{j}(15I-A_{S}Z_{j}(7I-A_{S}Z_{j})))$ (7)
@@
LaFiAttention approximates $A^{+}_{S}$ by $Z^{*}$ with Lemma 1. Following the empirical choice from Xiong et al. (2021), we run 6 iterations in order to achieve a good approximation of the pseudoinverse. Then, the softmax matrix $S$ used in self-attention is approximated by
$\hat{S}=softmax(\frac{Q\widetilde{K}^{T}}{\sqrt{d_{q}}})Z^{*}softmax(\frac{\widetilde{Q}K^{T}}{\sqrt{d_{q}}}).$ (8)
@@ 4 Approach
Figure 1: The LaFiCMIL framework processes large files for classification. Initially, document chunks are transformed into embedding vectors using BERT. A learnable category vector is then concatenated to these embeddings to form an augmented bag $X_{i}^{0}$ with $n^{\prime}=n+1$ instances. The LaFiAttention layer captures the inter-instance correlations within $X_{i}^{0}$. Operations within this layer, such as matrix multiplication ($\times$) and addition ($+$), are specified alongside the variable names and matrix dimensions. Key processes include sMEANS for landmark selections similar to Shen et al. (2018), pINV for pseudoinverse approximation, and DConv for depth-wise convolution. Classification is completed by passing the learned category vector through a fully connected layer.
@@ 4.1 Correlated Multiple Instance Learning
In the basic MIL concept Ilse et al. (2018), instances in a bag are independent and unordered. However, correlations may exist among instances within a bag, where the presence or absence of a positive instance can be influenced by other instances. In fact, when formulating large file classification as a MIL problem, correlations among instances can be found due to the presence of semantic dependencies between paragraphs. According to Theorem 2, these correlations can be exploited to reduce uncertainty in prediction. In other words, this relationship can be leveraged as additional information to boost the performance of long document classification tasks. We provide the mathematical definition of the Correlated Multiple Instance Learning (c-MIL) below.
c-MIL Formulation
Here, we consider a binary classification task of c-MIL as an example. Given a bag (i.e., a large file) $X_{i}$ composed of instances (i.e., small chunks) $\{x_{i,1},x_{i,2},...,x_{i,n}\}$, for $i=1,...,N$, that exhibit dependency or ordering among each other. The bag-level label is $Y_{i}$, yet the instance-level labels $\{y_{i,1},y_{i,2},...,y_{i,n}\}$ are not accessible. Then, a binary classification of c-MIL can be defined as:
$Y_{i}=\left\{\begin{array}[]{l}0,\indent if\sum y_{i,j}=0\indent y_{i,j}\in\{0,1\},j=1,...,n\\ 1,\indent otherwise\end{array}\right.$ (9)
$\hat{Y}_{i}=S(X_{i}),$ (10)

### v4

@@ Abstract
Transfomer-based models have significantly advanced natural language processing, in particular the performance in text classification tasks. Nevertheless, these models face challenges in processing large files, primarily due to their input constraints, which are generally restricted to hundreds or thousands of tokens. Attempts to address this issue in existing models usually consist in extracting only a fraction of the essential information from lengthy inputs, while often incurring high computational costs due to their complex architectures. In this work, we address the challenge of classifying large files from the perspective of correlated multiple instance learning. We introduce LaFiCMIL, a method specifically designed for large file classification. It is optimized for efficient training on a single GPU, making it a versatile solution for binary, multi-class, and multi-label classification tasks. We conducted extensive experiments using seven diverse and comprehensive benchmark datasets to assess LaFiCMIL’s effectiveness. By integrating BERT for feature extraction, LaFiCMIL demonstrates exceptional performance, setting new benchmarks across all datasets. A notable achievement of our approach is its ability to scale BERT to handle nearly 20 0002000020\,00020 000 tokens while training on a single GPU with 32GB of memory. This efficiency, coupled with its state-of-the-art performance, highlights LaFiCMIL’s potential as a groundbreaking approach in the field of large file classification.
Keywords:
Large file classification Multiple instance learning
@@ 1 Introduction
Text classification is a fundamental task in Natural Language Processing (NLP), entailing the assignment of suitable label(s) to specific input texts [24, 34]. This process is crucial across various domains, including sentiment analysis [12], fake news detection [25], and offensive language identification [37], among others. Recent years have seen the emergence of attention-based models like Transformer [47], GPT [35, 36], and the BERT family [13, 16, 45], which have established state-of-the-art benchmarks in text classification tasks. However, the challenge of processing very long sequences remains a significant obstacle, mainly due to their high computational requirements when facing extremely huge number of tokens.
There are mainly two types of solutions in the literature to address long token sequences: ① extending the input length limit by employing a sliding window to attention, such as Longformer [7], and ② dividing long documents into segments and recurrently processing the Transformer-based segment representations, such as RMT [10, 9]. Nevertheless, Longformer inherently struggles with global context capture. The sliding window mechanism can lead to a fragmented understanding of the overall sequence, as it primarily focuses on local context within each window. This becomes particularly challenging when dependencies span beyond the scope of these localized windows, a common occurrence in complex or extremely long text sequences. Similarly, the recurrent processing of RMT can also lead to information loss, especially when dependencies or context need to be carried over long sequences of text. Each recurrent step has the potential to dilute or overlook critical information from previous segments, leading to a gradual decay in context retention as the sequence progresses.
Recently, significant progress in large language models (LLMs) like GPT-4 [1], ChatGLM [52], and Llama 2 [46] has showcased their impressive capabilities in various NLP tasks. Despite these advances, directly applying large language models (LLMs) to text classification through prompt engineering has not achieved optimal performance, and lightweight Transformer models, such as RoBERTa [29, 55], continue to excel in this essential task. This is demonstrated by RGPT [55], an adaptive boosting framework designed to fine-tune LLMs for text classification. Although RGPT sets new benchmarks, it has only been validated on short texts and demands substantial computational resources, requiring eight A100-SXM4-40GB GPUs, and a total of 320 GB of memory. Consequently, developing a more resource-efficient and effective strategy for classifying long documents remains a critical challenge.
@@
In this work, we innovatively tackle this challenge by adopting Multiple Instance Learning [41, 53] (MIL), wherein we conceptualize a large file as a ’bag’ and its constituent chunks as ’instances’ within the MIL framework. We introduce LaFiCMIL, a simple yet effective Large File Classification approach based on correlated Multiple Instance Learning. On the one hand, as proven in Theorem 1 (cf., Section 3), a MIL score function for a bag classification task can be approximated by a series of sub-functions of the instances. This inspires us to split a large file into smaller chunks and extract their features separately using BERT. On the other hand, we aim to guide the model to learn high-level overall features from all instances, rather than deriving the final bag prediction from instance predictions based on a simplistic learned projection matrix. In addition, in contrast to the basic version of MIL [19], where instances within the same bag exhibit neither dependency nor ordering among one another, we claim that the small chunks from the same large file are correlated in some way (e.g., semantic dependencies in paragraphs). This implies that the presence or absence of a positive instance in a bag can be influenced by the other instances contained within the same bag. As a result, relying on our computationally efficient LaFiAttention layer, our approach is capable of efficiently extracting correlations among all chunks as additional information to boost classification performance.
@@
  • The training of LaFiCMIL is super efficient, which requires only 1.86$\times$ training time than the original BERT, but is able to handle 39$\times$ longer sequence on a single GPU with only 32 GB of memory.
@@
  • We share the datasets and source code to the community at: https://github.com/Trustworthy-Software/LaFiCMIL
@@ 2.1 Large File Classification
In recent years, significant efforts have been made to alleviate the input limit of Transformer-based models to handle large files. One notable example is Longformer [7], which extends the limit to $4096$ tokens using a sparse attention mechanism [51]. CogLTX [14] chooses to identify key sentences through a trained judge model. Alternatively, ToBERT [32] and RMT [10, 9] segment long documents into fragments and then aggregate or recurrently process their BERT-based representations. Recently, two simple BERT-based methods proposed in [33] achieved state-of-the-art performance on several datasets for long document classification. Specifically, BERT+Random selects random sentences up to 512 tokens to augment the first 512 tokens. BERT+TextRank augments the first 512 tokens with a second set of 512 tokens obtained via TextRank [31]. They also provide a comprehensive evaluation to compare the relative efficacy of various baselines on diverse datasets, which revealed that no single approach consistently outperforms others across all six benchmark datasets, encompassing different classification tasks such as binary [22], multi-class [26], and multi-label classification [6, 11].
@@ 2.2 Multiple Instance Learning
Multiple Instance Learning (MIL) has attracted increasing research interests and applications in recent years. The application scenarios of MIL span across various domains [20, 44, 18], but the most prominent one is Medical Imaging and Diagnosis. Particularly, there has been a growing trend towards developing MIL algorithms for medical whole slide image analysis [21, 50]. MIL approaches are broadly divided into two categories. The first group makes bag predictions based on individual instance predictions, typically using average or maximum pooling methods [15, 27, 30, 42]. The second group aggregates instance features to form a high-level bag representation, which is then used for bag-level predictions [28, 41, 53]. Although instance-level pooling is straightforward, aggregating features for bag representation has proven more effective [48, 41].
A fundamental assumption behind Multiple Instance Learning (MIL) is that instances within a bag are independent, but this might not always be true in real world. To tackle this, some research turns to Correlated Multiple Instance Learning (c-MIL), which assumes instances in a bag are correlated [57, 54, 41]. This approach recognizes that instances in a bag can influence each other. However, using c-MIL for large file classification remains under-explored.
@@ 3 Technical Preliminaries
In this section, we describe several essential technical preliminaries that inspire and underpin the design of LaFiCMIL. We first present a pair of theorems that are the fundamental principles of c-MIL.
Theorem 1
Suppose $S:\chi\rightarrow\mathbb{R}$ is a continuous set function w.r.t Hausdorff distance [39] $d_{H}(.,.)$. $\forall\varepsilon>0$, for any invertible map $P:\chi\rightarrow\mathbb{R}^{n}$, $\exists$ function $\sigma$ and $g$, such that for any set $X\in\chi$:
$\small|S(X)-g(P_{X\in\chi}\{\sigma(x):x\in X\})|<\varepsilon$ (1)
The proof of Theorem 1 can be found in [41]. This theorem shows that a Hausdorff continuous set function $S(X)$ can be arbitrarily approximated by a function in the form $g(P_{X\in\chi}\{\sigma(x):x\in X\})$. This insight can be applied to MIL, as the mathematical definition of sets in the theorem is equivalent to that of bags in MIL framework. Consequently, the theorem provides a foundation for approximating bag-level predictions in MIL using instance-level features.
Theorem 2
@@
$\small\begin{split}H(\theta_{1},\theta_{2},...,\theta_{n})&=\sum_{t=2}^{n}H(\theta_{t}|\theta_{1},\theta_{2},...,\theta_{t-1})+H(\theta_{1})\\ &\leq\sum_{t=1}^{n}H(\theta_{t})\end{split}$ (2)
The proof of Theorem 2 can be found in [41]. This theorem demonstrates that the information entropy of a bag under the correlation assumption is smaller than the information entropy of a bag under the i.i.d. assumption. The lower information entropy in c-MIL suggests reduced uncertainty and the potential to provide more information for bag classification tasks. In Section 4.1, we introduce c-MIL, and in Section 4.2, we derive the efficient LaFiCMIL based on c-MIL.
Next, we present the necessary preliminaries for our efficient attention layer inspired by the Nyströmformer [49], referred as LaFiAttention, which performs as a sub-function within our LaFiCMIL. In the original Transformer [47], an input sequence of $n$ tokens of dimensions $d$, $X\in{\bf R}^{n\times d}$, is projected using three matrices $W_{Q}\in{\bf R}^{n\times d_{q}}$, $W_{K}\in{\bf R}^{n\times d_{k}}$, and $W_{V}\in{\bf R}^{n\times d_{v}}$, referred as query, key, and value respectively with $d_{k}=d_{q}$. The outputs $Q,K,V$ are calculated as
$\small Q=XW_{Q},\leavevmode\nobreak\ K=XW_{K},\leavevmode\nobreak\ V=XW_{V}$ (3)
@@
$\small D(Q,K,V)=SV=softmax(\frac{QK^{T}}{\sqrt{d_{q}}})V$ (4)
@@
$\small S=softmax(\frac{QK^{T}}{\sqrt{d_{q}}})=\begin{bmatrix}A_{S}&B_{S}\\ F_{S}&C_{S}\end{bmatrix}$ (5)
@@
$\small\hat{S}=softmax(\frac{Q\widetilde{K}^{T}}{\sqrt{d_{q}}})A^{+}_{S}softmax(\frac{\widetilde{Q}K^{T}}{\sqrt{d_{q}}}),$ (6)
where $\widetilde{Q}=[\widetilde{q_{1}};...;\widetilde{q_{m}}]\in{\bf R}^{m\times d_{q}}$ and $\widetilde{K}=[\widetilde{k_{1}};...;\widetilde{k_{m}}]\in{\bf R}^{m\times d_{q}}$ are the selected landmarks for inputs $Q=[q_{1};...;q_{n}]$ and $K=[k_{1};...;k_{n}]$, $A^{+}_{S}$ is the Moore-Penrose inverse [8] of $A_{S}$.
Lemma 1
For $A_{S}\in{\bf R}^{m\times m}$, the sequence $\{Z_{j}\}^{j=\infty}_{j=0}$ generated by [38],
$\small Z_{j+1}=\frac{1}{4}Z_{j}(13I-A_{S}Z_{j}(15I-A_{S}Z_{j}(7I-A_{S}Z_{j})))$ (7)
@@
LaFiAttention approximates $A^{+}_{S}$ by $Z^{*}$ with Lemma 1. Following the empirical choice from [49], we run 6 iterations to achieve a good approximation of the pseudoinverse. Then, the softmax matrix $S$ is approximated by
$\small\hat{S}=softmax(\frac{Q\widetilde{K}^{T}}{\sqrt{d_{q}}})Z^{*}softmax(\frac{\widetilde{Q}K^{T}}{\sqrt{d_{q}}}).$ (8)
@@ 4 Approach
Figure 1: LaFiCMIL. Initially, document chunks are transformed into embedding vectors using BERT. A learnable category vector is then concatenated to these embeddings to form an augmented bag $X_{i}^{0}$ with $n^{\prime}=n+1$ instances. The LaFiAttention layer captures the inter-instance correlations within $X_{i}^{0}$. Operations within this layer, such as matrix multiplication ($\times$) and addition ($+$), are specified alongside the variable names and matrix dimensions. Key processes include sMEANS for landmark selections similar to [43], pINV for pseudoinverse approximation, and DConv for depth-wise convolution. Classification is completed by passing the learned category vector through a fully connected layer.
@@ 4.1 Correlated Multiple Instance Learning
In the basic MIL concept [19], instances in a bag are independent and unordered. However, correlations among chunks of a large file exist due to the presence of semantic dependencies between paragraphs. According to Theorem 2, these correlations can be exploited to reduce uncertainty in prediction. In other words, this relationship can be leveraged as additional information to boost the performance of long document classification tasks. The Correlated Multiple Instance Learning (c-MIL) is defined as below.
Here, we consider a binary classification task of c-MIL as an example. Given a bag (i.e., a large file) $X_{i}$ composed of instances (i.e., chunks) $\{x_{i,1},x_{i,2},...,x_{i,n}\}$, for $i=1,...,N$, that exhibit dependency or ordering among each other. The bag-level label is $Y_{i}$, yet the instance-level labels $\{y_{i,1},y_{i,2},...,y_{i,n}\}$ are not accessible. Then, a binary classification of c-MIL can be defined as:
$\small Y_{i}=\left\{\begin{array}[]{l}0,\indent if\sum y_{i,j}=0\indent y_{i,j}\in\{0,1\},j=1,...,n\\ 1,\indent otherwise\end{array}\right.$ (9)
$\small\hat{Y}_{i}=S(X_{i}),$ (10)